OpenStudy (anonymous):

I'm stuck on this one))Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 5 a.x = (x – 6) + 5 b.(y + 5) + y = 6 c.x + (y + 5) = 6 d.x + y = y + 5

pooja195 (pooja195):

Well we did a bunch of these so im assuming you learned about the substitution method We already know what x='s its given x + y = 6 <---e1 x = y + 5 <---e2 What do you get when you substitute e2 into e1?

OpenStudy (anonymous):

I don't know..I'm confused

pooja195 (pooja195):

Ok they told us x=y+5 correct?

OpenStudy (anonymous):

Yes

pooja195 (pooja195):

Ok we know the value of x its y+5 If you replace the x in equation 1 with y+5 what do we get?

OpenStudy (anonymous):

1= y + 5

pooja195 (pooja195):

Where did you get 1 from?

OpenStudy (anonymous):

Sorry I misread it.. I thought you said replace x with 1..

OpenStudy (anonymous):

Is it D?

pooja195 (pooja195):

This is equation 1 \(\huge\rm\color{red}{x + y = 6}\) \(\huge\rm\color{green}{x=y+5}\) If you take: \(\huge\rm\color{green}{y+5}\) And replace x in: \(\huge\rm\color{red}{x+ y = 6}\) With \(\huge\rm\color{green}{y+5}\) What do you get?

pooja195 (pooja195):

D is incorrect

OpenStudy (anonymous):

Give me a sec to solve it

pooja195 (pooja195):

Dont solve it just plug it in

pooja195 (pooja195):